OpenStudy (anonymous):

Help with this difference quotient? Let f(x)= 7-5x-2x^2 Use the difference quotient: f(x+h)-f(x)/h I'm just not sure how to apply the f(x) in the f(x+h).

Nnesha (nnesha):

first replace x by ( x+h) in f(x)

Nnesha (nnesha):

\[\huge\rm f(\color{Green}{x+h})=-2(\color{green}{x+h})^2 -5(\color{green}{x+h}) +7\] like this

Nnesha (nnesha):

can you solve that ?? :-)

OpenStudy (anonymous):

Oh okay so they would separated. That makes sense.

Nnesha (nnesha):

yeah replace x by x+h and that's just f(x+h) part of the formula

OpenStudy (anonymous):

((-2(x+h)^2)-(5x+h)+7)-(7-5x-2x^2)/h Then I'll simplify

Nnesha (nnesha):

yep right :-)

OpenStudy (anonymous):

((-2x-2h)^2)-(5x+5h)+7)-(7-5x-2x^2)/h (-2x-2h)-(5x+5h)-(-5x-2x^2)/h (2x^2-2x-7h)/h

Nnesha (nnesha):

\[\large\rm\frac{ (-2\color{red}{{(x+h})^2} -5(\color{green}{x+h}) +7 )-(7-5x-2x^2) }{ h }\] first sove parenthses you can't distribute by 2 first first step solve parenthses (x+h)^2foil

Nnesha (nnesha):

solve*

OpenStudy (anonymous):

Oh pellet duh didn't see that, I should be writing it down then solving via computer aha

Nnesha (nnesha):

hmm parentheses first ! try again o^_^o

OpenStudy (anonymous):

Okay, I don't know if what I got it right but 2(2x^2+2xh+h^2)-5h/h

Nnesha (nnesha):

okay (x+h)^2 = ??

OpenStudy (anonymous):

(2x^2+2xh+h^2)

Nnesha (nnesha):

2x^2 ??

OpenStudy (anonymous):

Whoops, I meant x^2

Nnesha (nnesha):